satellite73 (satellite73):

one more from the complex plane. expand \[\frac{z^2-1}{(z^2+1)^2}\] about \[z_0=i\]

OpenStudy (anonymous):

i have the answer, just not sure of the algebra strategy. partial fractions?

OpenStudy (jamesj):

Like before, pull out the 1/(z-i) factor. Then do some long division to simplify the remaining expression.

OpenStudy (anonymous):

as in \[\frac{1}{z-i}\frac{z^2-1}{(z-i)(z+i)^2}\]?

OpenStudy (jamesj):

Pull out (z-i)^2.

OpenStudy (anonymous):

\[\frac{1}{(z-i)^2}\frac{z^2-1}{(z+i)^2}\]

OpenStudy (jamesj):

Yes. Now, get rid of the z^2 in the numerator and you'll have a cz + d expression for some complex numbers c, d.

OpenStudy (anonymous):

ok thanks again. looks like this is just going to be some annoying algebra

OpenStudy (anonymous):

i'd rather compute the limit of a reimann sum

OpenStudy (jamesj):

lol. You've seen that? This strobe guy is real retrice

OpenStudy (jamesj):

restrice? no. a-s-s.

OpenStudy (anonymous):

yeah i got a chuckle from that one

OpenStudy (anonymous):

there must be something about math (maybe the fact that one is usually right or wrong) that turns people into opinionated retrices pieces

OpenStudy (pokemon23):

chess time

OpenStudy (anonymous):

hmmm this is not working the way i thought it would. maybe i made a mistake

OpenStudy (anonymous):

e4

OpenStudy (pokemon23):

e5

OpenStudy (anonymous):

n f3

OpenStudy (pokemon23):

hmmm